OpenStudy (anonymous):

Help! cube root of x^2-9x = -2

OpenStudy (anonymous):

Is (x^2-9x) all under the cube root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2-9*x-(-2)=0

OpenStudy (anonymous):

x =(9-√73)/2= 0.228 2. x =(9+√73)/2= 8.772

Nnesha (nnesha):

ok so your question is  is this right

Nnesha (nnesha):

let me know please @gabz12 thanks :)

OpenStudy (anonymous):

Can you just raise both sides to the third power?

OpenStudy (anonymous):

yes @Nnesha

Nnesha (nnesha):

 if this your question you have to take cube both side so left side cubic cancel out